40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 25 of 53  Not logged in ELOG logo
ID Date Author Type Category Subject
  1410   Thu Mar 20 21:54:10 2014 EvanNotesNoiseBudgetextracting phiL phiH

Quote:

 

phiL x10^-4 phiH x10^-4
0.1 9.6
0.5 9.05
1 8.33
1.5 7.61
2 6.89
3 5.35
 

 

 

 I independently computed the Hong result using the same assumptions (bulk and shear loss angles are equal, and no light penetration). I find

phiL x10^-4 phiH x10^-4
0.1 9.6
0.5(3) 9.0(6)
1.0(2) 8.3(5)
1.5 7.6
2.0 6.9
3.0 5.5

where I have included uncertainties for the Penn and Crooks measurements.

  1409   Thu Mar 20 18:52:40 2014 EvanDailyProgressLaserCTN north cavity locking

Laser is locked to north cavity, with slow PID loop engaged.

Current north laser slow DC voltage: 6.55 V, with some slow upward drift

TTFSS settings: 634 fast, 888 common (very lucky!)

 

  1408   Thu Mar 20 18:09:33 2014 taraNotesNoiseBudgetextracting phiL phiH

 Since we measured thermal noise from the coating(QWL, SiO2/Ta2O5), we want to extract loss angles of each materials.  The losses are about a factor of 2 higher than the numbers reported in the literature.

So far, there are 3 calculations we have been using for coating noise estimation.

  • Nakagawa , this calculation assumes Young’s moduli , Poisson ratio of the coating and the substrate are the same. The coating is s a single thin layer.  Sx ~ phiC . If we fit this calculation to our result, the coating loss is about 4.15e-4. This number agrees with the result from Numata2003, and agree with both our measurement from short and long cavities.
  • Harry2002:  Sx is function of ( phi_perp, phi_para), both phi_perp and phi_para are functions of( Y,Yc,sigma,sumac, phiL, phiH)
  • Hong2013,  Sx is function of phi_bulk and phi_shear. If we assume that phi bulk and shear are the same for each material phi bulk L = phi shear L = phi L, Sx ~ const1 phi L  + const2 * phi H,,  

In essence, both Harry's and Hong's result can be written as a linear combination of phiL and phiH.  I used Harry result to compare with Hong to see if there is any differences in the result or not, but both gave me the same answer.

==calculation==

The calculation is attached below. I made sure that the calculation from Hong and Harry are correct by choosing the elastic properties of the coatings to be the same as that of substrate and checking the the results agree with Nakagawa's. So the code should be correct.

Then, I varied phiL and phiH to match the measurement. The measurement is represented by the prediction by Nakagawa with the fitted loss (phiC = 4.15e-4).

Both calculations gave the similar relation between phiH (phi tantala) and phiL (phi silica) to match the measurement: 

 phiH = -1.4 phiL + 9.7  (Hong)

 phiH = -1.44 phi L + 9.77 (Harry) (assume sigma1 = sigma 2 = 0) 

The problem is if we use the nominal numbers from various reports, phiL ~ 1e-4, phiH ~ 4e-4.  The result will be off by almost a factor of 2. For example, for phiL = 1e-4, this means phi H has to be 8.4e-4.  Or if phi H is chosen to be 4e-4, phi L will be ~4e-4 as well. It seems that our result is higher than the predictions (under some assumptions).

 

Table1 below shows some possible values of phiL and phiH extracted from our result and the calculation.

phiL x10^-4 phiH x10^-4
0.1 9.6
0.5 9.05
1 8.33
1.5 7.61
2 6.89
3 5.35
 

 But we have good evidences from Numata and short/long cavities (spot size dependent) to believe that the measurement is real coating thermal noise . The reason why the prediction is smaller than the measurement could be that the losses is actually higher in our coating. Most ring down measurements were done after 2002 while our coatings were fabricated around 1997. Coating vendors might become more careful about loss and improved their process. But the result from Numata was out in 2003, and it is about the same as ours, so I'm really not sure what can we say about this.

==numbers from literature==

Penn2003: (disc ring down)  phiL = (0.5 +/- 0.3) x10^-4 ,  phiH = (4.4 +/- 0.2) x10^-4

Numata2003 (direct measurement)    phiC = 4.4e-4;

Crooks2004(disc ring down)    phiL = 0.4+/-0.3  x10^-4,   phiH = (4.2+/-0.4)x10^-4   (the frequency dependent part is ignored)

Crooks2006: (disc ring down)   phiL = (1.0 +/- 0.2) x10^-4     phiH = (3.8 +/- 0.2) x10^-4    (small change in TE calculation from previous paper)

Martin2009 : (blade)                                                      phiH = 3+/- 0.5 x10^-4  (at 300K)

Martin2010: (blade)                                                       phiH = (2.5-5) x10^-4  ( heat treated at 600C, several frequencies)

 LMA2014: (blade)   phiL = (0.43+/-0.02) x10^-4        phiH = (2.28 +/- 0.2) x10^-4 

 

 

Attachment 1: Hong_coat_BR.zip
  1407   Tue Mar 4 18:53:48 2014 taraNotesRefCavcomsol simulation for cavity suspension

I tried to estimate the coupling from seismic to displacement noise. With the common mode rejection taken into account, the coupling from vertical acceleration to differential length between the two cavities is about 6x10^-12.

 

  • With random errors add in the position (+/- 1mm) of the support position, I used COMSOL to find out the coupling from acceleration to strain noise to a single cavity. Note that I did not use Gaussian distribution because sometimes the support position will be place away from the cavity. so I used random error with confined limit to +/- 1mm.
  • Then I histogram it. The result is shown in the figure below.

hist_small_mount_err.png 

  • Take some common mode rejection into account. Since most of the time, the coupling will fall between 0.98 x10^-10 to 1.04 x10^-10  [s^2/m].  We can assume an upperbound that one cavity has the coupling of 1.04x10^-10 and another has 0.98x10^-10. Thus, the effective coupling for the differential strain is (1.04-0.98)x10^-10 = 6x10^-12. 
  • I have to admit that the result is very weird. With the ideal support position, the coupling for a single cavity is also about 6x10^-12. I did not expect that by changing the positions slightly, the coupling for each cavity can be worse by 200 times. I suspect that COMSOL might add some errors as well, see psl:1056. The cavity does not bend down symmetrically, even with the perfectly symmetric support position, but I did not think it will be this huge.
Attachment 2: hist_small_mount_err.fig
  1406   Sun Feb 9 20:37:38 2014 EvanSummaryNoiseBudgetNoise budget fitting: result

Quote:

Quote:

In order to get the systematic uncertainty on ϕc, we need uncertainties in other parameters that enter the noise budget. Specifically:

  • Spot size, w. Currently using (182.0 ± 0.4) µm. The nominal comes from eq. 47 in Kogelnik and Li. This uncertainty can be propagated forward uncertainties in the following:
    • Cavity length, L. The spacer drawing (CTNLab/drawings/mechanical_drawings/dual_refcav/cavity_spacer_1.45inx1.5in.PDF) gives the length as (1.45 ± 0.01) inches.
    • Mirror radius, R. Nominal is 0.5 m; no idea about the uncertainty. It is not given in the test document. Currently using (500 ± 3) mm; i.e., an 0.5% uncertainty. This is the uncertainty claimed in the CVI catalog.
  • Coating thickness, d. Nominal is 4.5 µm; no idea about the uncertainty. Currently using (4.53 ± 0.07) µm from Tara's calculation (given in the reply to this elog post).
  • Substrate elastic modulus, Es. Using (73.1 ± 0.1) GPa as estimated from figure 29 in McSkimin 1953 (doi 10.1063/1.1721449), which is a (heavily cited) reference I found on the NIST ceramics database.
  • Substrate Poisson ratio, σs. Using 0.170 ± 0.004 as estimated from figures 29 and 30 in McSkimin (for Young and shear moduli, respectively) and then propagating the error forward to the Poisson ratio.
  • Cavity temperature. Should use 306(1) K.
  • Measurement uncertainty. Given an estimated PSD S(f) that is obtained by averaging M FFTs, the uncertainty should be S(f)/M1/2. Tara says to take M = 50. Possibly need to adjust formula to account for Welch overlap.

The above uncertainties are enough to estimate the statistical and systematic uncertainties on a fit to ϕc using the Nakagawa/Harry formula for a thin, lossy coating. By minimizing an appropriately weighted chi-squared function from 50 Hz to 500 Hz and then taking into account the above substrate and coating uncertainties, I find ϕc = (4.15 ± 0.03 stat ± 0.08 sys) × 10−4. More details will follow, and there may need to be some refinement (e.g., I still haven't dealt with the Welch overlap issue).

This has required adjusting the values of the Young modulus and Poisson ratio from their previous values (72.7 GPa and 0.167, respectively). I haven't checked these changes into the SVN.

On Tara's suggestion, I've done a fit to a coating loss angle with a power-law frequency dependence. The results are highly dependent on the band chosen for the fit (see attached plot).

For comparison, for a fit to a frequency-independent loss angle, the dependence on the band is much less prominent. For 50 Hz to 200 Hz, I get 4.12(3) × 10−4, and for 50 Hz to 700 Hz, I get 4.21(3) × 10−4.

Attachment 1: lossfit.pdf
lossfit.pdf
  1405   Tue Feb 4 01:11:01 2014 taraNotesRefCavcomsol simulation for cavity suspension

 I'm trying to estimate the coupling between seismic to displacement noise of the cavity using COMSOL. 

From the design, the strain due to the seismic noise is about 2e-11.  But we want to see what happen if the support positions are moved away from the specified points a bit. This time the model is a whole cavity, not just 1/8 as I did before. This is to see results of the mis-positioned support points. However, COMSOL has some problems

  • The solution for the full sized cavity is not symmetric even with the optimum support points. I think this is because the mesh size is not physically symmetric, but the result gets better (the sagging becomes more symmetric) when the mesh size is smaller. However, the calculation time is quite long (~3.5 minutes for each run).
  •  I cannot adjust the angle, I can only adjust the beam line position. I'm not sure why COMSOL has this problem. It is not a syntax problem. I ran the iteration and it gave me two results , then the error message poped out. So I only changed the support positions along the beam line direction 
  • The random I used is white, where the error is +/- 0.5 mm. I could not use Gaussian, because sometimes the support positions were out of the spacer. I think It should be fine for now to see how the result will be

Right now I have ~30 data points after 4 hrs of running the simulation. I'll get a bit more data and will see how it goes when I histogram it.

  1404   Tue Feb 4 00:31:32 2014 EvanSummaryNoiseBudgetNoise budget fitting: result

Quote:

Quote:

In order to get the systematic uncertainty on ϕc, we need uncertainties in other parameters that enter the noise budget. Specifically:

  • Spot size, w. Currently using (182.0 ± 0.4) µm. The nominal comes from eq. 47 in Kogelnik and Li. This uncertainty can be propagated forward uncertainties in the following:
    • Cavity length, L. The spacer drawing (CTNLab/drawings/mechanical_drawings/dual_refcav/cavity_spacer_1.45inx1.5in.PDF) gives the length as (1.45 ± 0.01) inches.
    • Mirror radius, R. Nominal is 0.5 m; no idea about the uncertainty. It is not given in the test document. Currently using (500 ± 3) mm; i.e., an 0.5% uncertainty. This is the uncertainty claimed in the CVI catalog.
  • Coating thickness, d. Nominal is 4.5 µm; no idea about the uncertainty. Currently using (4.53 ± 0.07) µm from Tara's calculation (given in the reply to this elog post).
  • Substrate elastic modulus, Es. Using (73.1 ± 0.1) GPa as estimated from figure 29 in McSkimin 1953 (doi 10.1063/1.1721449), which is a (heavily cited) reference I found on the NIST ceramics database.
  • Substrate Poisson ratio, σs. Using 0.170 ± 0.004 as estimated from figures 29 and 30 in McSkimin (for Young and shear moduli, respectively) and then propagating the error forward to the Poisson ratio.
  • Cavity temperature. Should use 306(1) K.
  • Measurement uncertainty. Given an estimated PSD S(f) that is obtained by averaging M FFTs, the uncertainty should be S(f)/M1/2. Tara says to take M = 50. Possibly need to adjust formula to account for Welch overlap.

The above uncertainties are enough to estimate the statistical and systematic uncertainties on a fit to ϕc using the Nakagawa/Harry formula for a thin, lossy coating. By minimizing an appropriately weighted chi-squared function from 50 Hz to 500 Hz and then taking into account the above substrate and coating uncertainties, I find ϕc = (4.15 ± 0.03 stat ± 0.08 sys) × 10−4. More details will follow, and there may need to be some refinement (e.g., I still haven't dealt with the Welch overlap issue).

This has required adjusting the values of the Young modulus and Poisson ratio from their previous values (72.7 GPa and 0.167, respectively). I haven't checked these changes into the SVN.

I've tried harmonizing the Hong result (eq 94) with the Nakagawa/Harry formula, but the phi_tantala that I extract is about 9e-4, which is twice as high as previously reported values. I've spent some time hunting for a missing factor of two, but cannot find one.

  1403   Mon Feb 3 23:56:07 2014 taraPhotosopticpacking mirror

 I"m packing the mirrors so that they are ready to be shipped to G. Cole. The mirrors are packed properly, see picasa.

  1402   Fri Jan 31 00:44:47 2014 EvanSummaryNoiseBudgetNoise budget fitting: result

Quote:

In order to get the systematic uncertainty on ϕc, we need uncertainties in other parameters that enter the noise budget. Specifically:

  • Spot size, w. Currently using (182.0 ± 0.4) µm. The nominal comes from eq. 47 in Kogelnik and Li. This uncertainty can be propagated forward uncertainties in the following:
    • Cavity length, L. The spacer drawing (CTNLab/drawings/mechanical_drawings/dual_refcav/cavity_spacer_1.45inx1.5in.PDF) gives the length as (1.45 ± 0.01) inches.
    • Mirror radius, R. Nominal is 0.5 m; no idea about the uncertainty. It is not given in the test document. Currently using (500 ± 3) mm; i.e., an 0.5% uncertainty. This is the uncertainty claimed in the CVI catalog.
  • Coating thickness, d. Nominal is 4.5 µm; no idea about the uncertainty. Currently using (4.53 ± 0.07) µm from Tara's calculation (given in the reply to this elog post).
  • Substrate elastic modulus, Es. Using (73.1 ± 0.1) GPa as estimated from figure 29 in McSkimin 1953 (doi 10.1063/1.1721449), which is a (heavily cited) reference I found on the NIST ceramics database.
  • Substrate Poisson ratio, σs. Using 0.170 ± 0.004 as estimated from figures 29 and 30 in McSkimin (for Young and shear moduli, respectively) and then propagating the error forward to the Poisson ratio.
  • Cavity temperature. Should use 306(1) K.
  • Measurement uncertainty. Given an estimated PSD S(f) that is obtained by averaging M FFTs, the uncertainty should be S(f)/M1/2. Tara says to take M = 50. Possibly need to adjust formula to account for Welch overlap.

The above uncertainties are enough to estimate the statistical and systematic uncertainties on a fit to ϕc using the Nakagawa/Harry formula for a thin, lossy coating. By minimizing an appropriately weighted chi-squared function from 50 Hz to 500 Hz and then taking into account the above substrate and coating uncertainties, I find ϕc = (4.15 ± 0.03 stat ± 0.08 sys) × 10−4. More details will follow, and there may need to be some refinement (e.g., I still haven't dealt with the Welch overlap issue).

This has required adjusting the values of the Young modulus and Poisson ratio from their previous values (72.7 GPa and 0.167, respectively). I haven't checked these changes into the SVN.

  1401   Wed Jan 29 21:36:53 2014 taraSummaryNoiseBudgeterror in spacer brownian noise

Spacer in BR noise

 

== COMSOL vs result from Kessler etal 2012==

     The analytical result from kessler2012, assume the force acts on whole surface of the spacer (with bore hole), I check this with COMSOL by comparing the result, similar to what I did in PSL:1075. The result agrees well within 2%. This verifies that COMSOL model is correct

 

==thermal noise level vs annulus thickness==


Typically, the contact surface between the spacer and the mirror is only a thin annulus, see psl:1199  . And the noise level is dependent on the actual area of contact. So I run the simulation to see the dependent of the stored energy (U) vs the annulus thickness. The annulus thickness is about 2 mm +/- 0.2 mm. The displacement noise is proportional to sqrt(U).

spacerBR_comsol.png

fig1: The stored energy as calculated by COMSOL, fitted with cubic polynomial.

The error from the contact area, the simulation result are small ~3% and 2%. These are smaller than the uncertainty of loss in bulk fused silica (can be from 10^-6 to 10^-7). The effect is still small in the total noise.

Attachment 2: spacerBR_comsol.fig
  1400   Wed Jan 29 05:44:51 2014 taraSummaryNoiseBudgetNoise budget fitting: need uncertainties

I looked into the uncertainty in coating thickness of the QWL SiO2/Ta2O5 coating The thickness of  4.53 +/- 0.07 um (~1.5%)seems to be appropriate.

The thermal noise level is directly proportional to the coating thickness, so we want to estimate its uncertainty. The error in the thickness is from

  • The uncertainties in nL and nH: since the physical thickness is lambda/(4*n), the error in n goes to the error in d.
  • Manufacturing process.

The errors in nL and nH are quite small, nL ~ 1.45 +/ 0.01, nH ~ 2.06+/- 0.01. (From the literature). I also looked around the error in IBS thickness control, they are usually better than 0.1 nm, IBS, but that is the current technology. In literature around 2000s, 2% error seems to be the number estimated for the thickness control (Sullivan 2000, Badoil 2007). As a quick check, I used the same assumption for error propagation similar to that of AlGaAs coating. The result gives ~ 4.53 +/- 0.07 um for coating thickness.

Note that the error here is smaller than the difference in coating thickness for the coatings with or without half wave cap.

For 28 Layer (with cap), the coating thickness is 4.53 um,  for 28 layer QWL, the coating thickness is 4.35 um. But after digging up all the information from REO, and peter king they agree that it is 28 QWL with half wave cap.  I tried to compare the calculation and the photothermal TF measurement, but the effect is too small to be conclusive about the structure. So the biggest error might come from the fact that the coating has cap or not. The error is about 4%.

  1399   Fri Jan 24 21:13:13 2014 taraDailyProgressDAQmDV in ATF

I'm trying to record beat measurement for a few days. The data will be taken from ATF using mDV. There are a few issues about mDV right now, I'm looking into it and asking around.

There is a problem with gps.m that converts the string to gps second. It is used in get_data where we specify the start time. I tried enter the gps second manually but it returns an empty time struct, and the get_data cannot be used.

 A reminder entry: psl:978

  1398   Wed Jan 22 16:18:37 2014 EvanSummaryNoiseBudgetNoise budget fitting: need uncertainties

In order to get the systematic uncertainty on ϕc, we need uncertainties in other parameters that enter the noise budget. Specifically:

  • Spot size, w. Currently using (182.0 ± 0.4) µm. The nominal comes from eq. 47 in Kogelnik and Li. This uncertainty can be propagated forward uncertainties in the following:
    • Cavity length, L. The spacer drawing (CTNLab/drawings/mechanical_drawings/dual_refcav/cavity_spacer_1.45inx1.5in.PDF) gives the length as (1.45 ± 0.01) inches.
    • Mirror radius, R. Nominal is 0.5 m; no idea about the uncertainty. It is not given in the test document. Currently using (500 ± 3) mm; i.e., an 0.5% uncertainty. This is the uncertainty claimed in the CVI catalog.
  • Coating thickness, d. Nominal is 4.5 µm; no idea about the uncertainty. Currently using (4.53 ± 0.07) µm from Tara's calculation (given in the reply to this elog post).
  • Substrate elastic modulus, Es. Using (73.1 ± 0.1) GPa as estimated from figure 29 in McSkimin 1953 (doi 10.1063/1.1721449), which is a (heavily cited) reference I found on the NIST ceramics database.
  • Substrate Poisson ratio, σs. Using 0.170 ± 0.004 as estimated from figures 29 and 30 in McSkimin (for Young and shear moduli, respectively) and then propagating the error forward to the Poisson ratio.
  • Cavity temperature. Should use 306(1) K.
  • Measurement uncertainty. Given an estimated PSD S(f) that is obtained by averaging M FFTs, the uncertainty should be S(f)/M1/2. Tara says to take M = 50. Need number of averages for each bin. Possibly need to adjust formula to account for Welch overlap.
  1397   Mon Jan 20 19:20:41 2014 EvanSummaryNoiseBudgetNoise budget fitting

I've added a χ2 minimization routine to nb_short_fit.m which looks for the value of ϕc (as defined by the Nakagawa formula) which makes the noise budget best fit the observed beat spectrum. For the weights in the χ 2 function, we need an estimate of the variance of the power in each bin. Ideally, we'd take multiple spectrum measurements and average them together. Since we only have a single measurement, for each bin I've taken the five bins on either side and computed the variance.

I performed the fit in the band from 26 Hz to 405 Hz because it looks like the total noise is dominated by coating Brownian noise in this region.

The first attachment shows χ2 as a function of ϕc. The routine assumes χ2 is parabolic in the neighborhood of the optimum value (which you can clearly see it is), and from this extracts the optimum value as well as the statistical uncertainty (which is given by the curvature of the parabola). From this the routine gives ϕc as 4.18(3) × 10−4, with a reduced of χ2 of 1.23.

From here, the next steps are

  • Settle on values for the lower and upper frequency limits for the fit. If the upper frequency is increased from 405 to 705 Hz, for example, the routine gives ϕc as 4.26(3) × 10−4, with a reduced of χ2 of 1.21. I think this is due (in part) to the fact that the noise budget is less than the measured noise near and above 1 kHz.
  • Perform the fit using the Harry formula, which includes the contributions from both loss angles rather than a single averaged value.
  • Compute the uncertainty arising from uncertainties in the other parameters. Tara has collected uncertainties for the material parameters in PSL:895, but we also need uncertainties in the spot size and the temperature.
Attachment 1: cost_func.pdf
cost_func.pdf
Attachment 2: nb_beat_short.pdf
nb_beat_short.pdf
Attachment 3: nb_beat_resid.pdf
nb_beat_resid.pdf
  1396   Fri Jan 17 12:46:29 2014 EvanSummaryNoiseBudgetNoise budget fitting: code fork

I've forked the noise budget code so that we can create a version that performs a fit to the coating loss angle. It is at CTNLab/simulations/noise_budget/iscmodeling/coating/iRefCav/nb_short_fit/nb_short_fit.m.

I've retooled the noise budget plot a bit. I've referred it to single-cavity length noise by multiplying the beat ASD by Lλ/(sqrt(2)c), where L = 3.7 cm. I've also combined some of the substrate noise, spacer noise, and technical noise traces so that there are not quite so many lines on a single plot. If we really want to display each trace individually, I think we should do so with a few separate plots (e.g., a thermal noise plot, a frequency/PLL noise plot, etc.). Fewer traces makes it easier for readers to make sense of the plot.

I'm going to start on writing the fitting code. For nonlinear least squares I'm used to using the Levenberg–Marquardt algorithm through scipy.optimize.curve_fit. I'll need to read up a bit on what's available in Matlab.

 

Attachment 1: nb_beat_short.pdf
nb_beat_short.pdf
  1395   Wed Jan 15 10:56:08 2014 EvanSummaryNoiseBudgetBeat measurement: residual

I've taken the total noise trace, interpolated it so that it uses the same frequency array as the measurement trace, and performed the quadrature subtraction of the two to get the residual. I've also converted the beat to single-cavity length noise by multiplying by Lλ/sqrt(2)c, with L = 3.7 cm.

Attachment 1: short_nb_residual.pdf
short_nb_residual.pdf
Attachment 2: short_nb_residual.fig
  1394   Thu Jan 9 09:09:44 2014 EvanDailyProgressRefCavSideband frequencies for refcavs

I've done (what I think is) more or less the same HOM computation as Tara for L = 3.7 cm and R = 0.5 m. Equation 51 in Kogelnik and Li gives the frequency of a mode with axial number q and transverse numbers m and n:

f / ffsr = q + 1 + (m + n + 1) arccos(1 − L / R) / π.

As a function of sideband frequency, I've plotted the detunings of the first 50 mode orders (and their sidebands) relative to the TEM00 carrier. Solid lines indicate carriers, and dashed lines indicate sidebands. The region from 32 to 35 MHz is right out, since the sidebands of mode orders 0 and 8 are very close.

I'm inclined to say that for R = 0.5 m alone, we should pick 26 MHz and 27 MHz, just because it's well out of the way of the forbidden 32 to 35 MHz region. As far as I know, the only other RF frequency to avoid is the PMC PDH frequency, which is 21.5 MHz.

Edit: I've done the above for L = 3.7 cm and R = 1 m, and the result is attached. If we want to accommodate R = 0.5 m and R = 1.0 m, it would be better to pick 36 and 37 MHz, or perhaps 23 and 24 MHz.

Attachment 1: pdh_choice_close.pdf
pdh_choice_close.pdf
Attachment 2: pdh_choice_close.pdf
pdh_choice_close.pdf
  1393   Tue Dec 31 19:33:47 2013 taraSummaryNoiseBudgetbeat measurement

I got a chance to measure beat measurement. The noise budget is updated and contains all dominant noise traces.

 

== Beat measurement ==

beat_2013_12_24.png

1) at DC to 10Hz, the contribution is mostly from RIN driven Photothermal noise and a bit of seismic noise, a small peaks around 10Hz is probably from the stack, not the cavity sagging. The hump from DC to ~ 50Hz disappear when it is quiet. I think it is mostly scattered light associated with the seismic noise, not displacement noise due to the vibration.

2) 10Hz to 1kHz is pretty much Coating Brownian noise.

3) At 1kHz and above, it is PLL readout noise and residual frequency noise from the laser, where the gain cannot suppress enough noise. This is mostly from ACAV. The residual frequency noise = free running noise / (1+ OLGTF). The measurement of the open loop gain is explained below.

 

==TTFSS Loop characterization==

The OLG TF of TTFSS is measured up to 10MHz and compared with the calculation. The schematic explaning how TTFSS actuates on the laser is shown below.

TTFSS.png

The freqeuncy discriminator can be measured from the slope of the error signal (from Common out1) while scanning the laser. For RCAV Dv = 1/ (194 kHz/V) and 1/(164kHz/V) for ACAV. with 1mW input power.

The adjustable gain stage can be tuned by turning the dial knob. At 400, gain=1, and the gain changes by 10dB with every 250click.

The PZT actuator has a gain of  4.5MHz/V (measured), and the EOM actuator is 15mRad/V (or 15mHz/f  Hz/V) (taken from the spec sheet).

OLG measurement is taken:  RCAV OLG is measured and plotted against the theoretical approximation, see the below figure.

RCAV_OLG.png

above: RCAV OLG TF. Note: The calculation and the measurement do not include the integrator with corner frequency at 4.6kHz.

 

There are some problems with ACAV loop and I could not increase the gain up as much as it used to be and the UGF is around only 200kHz , but the measurement matches the calculation. Right now RCAV servo has a better loop performance.

ACAV_OLGTF.png

 The calculated OLG TF trace(green) should go down at 1MHz or above because of the opamps' bandwidth. I used ideal Op Amps in the simulation because I don't have some op amps in my liso library. I'll see if I can fix it.

  1392   Wed Dec 18 21:05:28 2013 taraNotesoptic photothermal noise in AlGaAs: thickness resolution

We heard back from G. Cole about the thickness resolution in the AlGaAs coating manufacturing process will be around 0.5 A. So I'm checking how the noise budget will change by rounding up the physical thickness in opt V4 to the next 0.5A. The design will still work. The round up thickness is added in the google document (for opt v4 only).

The estimated growth rate of the crystal is 4.8A/s and shutter speed is assumed to have 0.1 sec time step. This means the smallest step of the thickness control is ~0.5A. So I round up the physical thickness to the next 0.5 A and calculate the coating properties.

1) Rounding up to the next 0.5 Angstrom. The truncating process acts like a random thickness variation in the optimized coatings with maximum error ~ 0.25 Angstrom. The averaged layer thickness is ~ 800 Angstrom.

 05Atrancate_err.png

 

2)Results when the layers physical thickness are round up to the closest 0.5 A. The noise budget does not change much.

05Atruncate_nb.png

05Atruncate_T.png

05Atruncate_err_ana.png

The coatings properties still hold, even with random error in parameters, thickness.

 

Note: For the error calculation I did before I used 1 sigma to be 1% for AlGaAs, and 0.5% for GaAs. The thinnest layer is AlGaAs at 35 A, so its sigma is about 0.35 A. The average thickness is 90 Angstrom, so the average error is about 0.9 A. The estimated error in the calibration process is already larger than the error from the truncation(0.25A). That's why the error analysis results are still valid.

Attachment 5: 05Atrancate_err.fig
Attachment 6: 05Atruncate_err_ana.fig
Attachment 7: 05Atruncate_nb.fig
Attachment 8: 05Atruncate_T.fig
  1391   Tue Dec 17 21:11:21 2013 ranaNotesElectronics EquipmentOLG of RCAV TTFSS

 

 This is a good first step, but there are two points:

1) You can NEVER use the SR785 for the FSS loop gain measurement since it doesn't go above the UGF. Use the RF network analyzer and do multiple sweeps to get the resolution using the Mott GPIB code.

2) This seems like its not good enough in the 10-100 kHz band.

3) We need to see how much phase is being lost because of the low modulation frequency and the high Q of the RFPD. Where is the plot comparing the servo model and the measurement?

  1390   Mon Dec 16 15:14:43 2013 taraNotesElectronics EquipmentOLG of RCAV TTFSS

 open loop gain transfer function of RCAV is measured. 

 1) how to measure OLG TF

  • see, PSL:592 about how to measure the OLG from TTFSS.
  • current schematic [add fig]

2) setup

  • 1mW input to the cavity
  • gain common/fast = 730/950
  • Mod depth =
  • Boost off 
  • U3 -> 11.84 dB

Result:

openloopTF.png

The requirement assumes that the residual frequency noise is 5% or less in the total noise. The servo performance is definitely ok for 1.45 inch cavity.

Attachment 2: openloopTF.fig
  1389   Mon Nov 25 15:28:12 2013 taraNotesRefCaveigenmodes of 1.45" refcav

I realized that we have not checked the eigenmodes of 1.45" cavity yet, so I used comsol to find out several modes. The lowest mode is ~ 46kHz, and the first longitudinal mode is about 60kHz. The frequencies are high enough so that the thermal noise calculation in dc- 10kHz frequency band can be done with quasi-static assumption.

 

1) I tried a simple cylindrical shape, with the dimension of the spacer. The result for the first longitudinal mode is 74KHz, the analytical result is ~ 77kHz, see PSL:1135. It seems that COMSOL's result and the analytical results are comparable.

spacer_only_z.png

 

2) Then I simulated the whole reference cavity. The lowest body mode is ~ 47kHz. The body expand-contract radially, and should not change the cavity beamline length that much. The first longitudinal mode is ~ 60kHz. The color on the surface shows the rms displacement from all direction.

 

spacer_br_8_edge.png

 

spacer_br_8_edge_z.png

Attachment 1: spacer_eigenmode.mph
  1388   Wed Nov 20 18:19:01 2013 taraDailyProgressNoiseBudgetphotothermal noise in SiO2/Ta2O5

I compared our beat measurement with results from Numata2003 and TNI. They agree well. I'm quite certain that we reach Brownian thermal noise from coatings.

 

 To make sure that what we measure is real Coating Brownian noise (It could be something else, i.e thermal noise in the support, spacer , or optical bond), we should compare our result to previous measurements to make sure that the numbers agree.

 Numata etal and TNI reported coating thermal noise measurement from suspended cavities (no spacer). They adjusted loss in the coatings to fit the measurement.  Phi coatings as reported in Numata is 4e-4 while TNI gives phi perp = phi_para = 2.7e-4.  Both agree with our result, see the plot below.  This means that our result is comparable with what they measured. It should be an evidence to support that we see real coating thermal noise, not contribution from something else (spacer, optical bond between the mirrors and the spacer).

beat_compare.png

Another evidence is from our previous measurement from 8" cavity.

  • The measurement also agrees with Numata's 2003 result, with phi coatings = 4e-4, see PSL:1018.
  • And the signal scales correctly with a factor of ~ 9 (from shorter cavity, and from smaller spotsize^2), seeT1200057. Had it been noise from optical bonding/ spacer (independent from spotsize), the scale factor would have been 8/1.45 ~ 5.5. The scale from substrate Brownian will also be different because of 1/w_spot dependent. Thermoelastic/ thermoopitc will have different slope.

So It is clear that our beat measurements from both 8" and 1.45" cavities are coating Brownian noise limited (around 50Hz-1kHz).

 

Attachment 2: beat_compare.fig
  1387   Tue Nov 12 15:27:32 2013 taraNotesDocumentationthesis on ctn

I created an svn folder for my thesis on CTN measurement.

It can be found here

 

 

  1386   Mon Nov 11 19:37:13 2013 taraDailyProgressopticredo- PMC path

PMC path is back, I aligned the polarization of the input beam to the BB EOM for TTFSS. The visibility of PMC is now ~ 80%.

  1385   Fri Nov 8 03:36:44 2013 taraDailyProgressopticredo- PMC path

I'm re-arranging the optics in PMC path a bit. The work is in progress, so ACAV path is still down.

I'm investigating why ACAV TTFSS performance is worse than that of RCAV. One thing is that ACAV has the PMC. This area has not been optimized for awhile, so I'm checking everything.

  1384   Thu Nov 7 05:08:13 2013 taraDailyProgressNoiseBudgetphotothermal noise in SiO2/Ta2O5

I add the photo thermal noise effect in the noise budget. With ISS, photothermal noise should be sufficiently small.

 

What I did

  • Measure beat
  • Measure RIN after ACAV and RCAV
  • Measure TF between TRANSPD and beat, compare the result with Farsi's calculation to determine the absorption (8ppm, with Finesse = 1e4) [add more details]
  • Apply the measured RIN to Farsi calculation to get the conversion from RIN to frequency noise ( I did not use the measured TF because I have not measured the whole range yet, and the calculation matches the measurement quite well).

beat_2013_11_07.png

Comment about the beat

  • At DC -30 Hz, the noise seems to be a combination of photothermal noise, and seismic induced scattered light. Air spring might not help as much as I thought.
  • Above 2kHz, it's not clear if it is gain limited on ACAV loop or not, but this is likely. We can check by measure the PSD of the error signal and convert it to frequency noise.
  • Frequency stabilization of ACAV is significantly inferior than that of RCAV. I don't know if it is the result from PMC or not. More investigation is needed.

Note about RIN measurement

  • RIN (measured behind the cavities) depends considerably on the TTFSS gain, luckily, at optimum gain level, RIN is pushed down enough.
  • RIN from ACAV is almost a factor of 10 worse than that of RCAV @ the optimum gain setting
  • There might be coupling from BB EOM to RIN (due to the mismatches E field between the EOM and the beam). This may explains why RIN is getting worse if common gain is increased a bit before the loop oscillate. Will check that.

 

Note about loss angles: For  SiO2 and Ta2O5 loss angles = 1e-4 and 7.5e-4 (a factor of 3 above the regular number), the noise budget matches the measurement well. I'll see if it is the same for the data from 8" cavities or not.

Attachment 2: beat_2013_11_07.fig
  1383   Wed Nov 6 01:14:58 2013 taraDailyProgressElectronics EquipmentTTFSS

We made a mistake by choosing the input power to the cavities to be 0.25 mW, so today I turned them back to 1mW and measure the beat.

beat_2013_11_05.png

Setup:

  • input power: 1mW
  • TTFSS gain (C/F): RCAV (760/980), ACAV (630/650)

Note about the measurement:

  • The noise at error point from ACAV is pretty high (~up to 100 nV/rtHz around 6 kHz). Better characterization will be done later to see if the suppression is enough or not. I made sure that this measurement is good up to ~2kHz, (this was done by changing the gain level a bit and beat level did not change).
  • Table was floated, but the air springs was not activated. I hope we will get better signal around 20-100 Hz once the air spring is re-installed.
  •  Intensity noise->photothermal around a few hundred mHz, cause the PLL to drift away from the input range. This effect becomes worse when the input power is increased.
  • RIN from ACAV is about a factor of 10 higher than that of RCAV.
  • I measured the beat with ISS on/off on ACAV, nothing were significantly different. So maybe it is not a problem for now.

To do next:

  • I'll compare this measurement with the one from 8" cavity to see if the results agree or not. They are different mirrors, but from the same coating run. I suspect that the loss in these mirrors are higher than what we estimate (Ta2O5=2.5e-4, SiO2 1e-4).
  • Think about the error in the measurement (calibration, spotsize etc).
Attachment 2: beat_2013_11_05.fig
  1382   Tue Nov 5 10:34:09 2013 EvanDailyProgressBEATNew beat measurement

[Tara, Evan]

Yesterday, we did a few final bits of optimization and then re-measured the beat spetrum.

Specifically:

  • Tara tweaked the north RFPD cable to symmetrize the north PDH error signal.
  • For each path, we hooked the RF output of each RFPD directly into the HP4395A and minimized the RFAM at 14.75 MHz (the PDH frequency).
    • For south, the initial RFAM was −107 dBm, and by adjusting the HWP before the resonant EOM, Tara was able to get it down to −112 dBm. This resulted in no visible change to the error signal offset (10 mV, compared to 300 mV peak-to-peak).
    • For north, the initial RFAM was −105 dBm, and by adjusting the HWP before the broadband EOM + resonant EOM (there is no intervening waveplate), Tara was able to get it below the noise floor of the spectrum analyzer (i.e., < −125 dBm). This resulted in a change to the error signal offset from 28 mV to 19 mV (compared to 300 mV peak-to-peak).
  • We measured the slopes of the error signals in order to get the calibration from voltage to frequency.
  • We locked the cavities with the slow digital control engaged. The south TTFSS gain was 712 slow and 796 fast, and the north TTFSS gain was 900 common and 900 fast. The south TTFSS control signal still has strong (~ 2 V peak-to-peak) oscillations at hundreds of kilohertz.
  • We measured the beat, the PDH error signals, and the RIN spectra. The modulation setting on the Marconi was 1 kHz, so the conversion factor is the usual 710 Hz/V.
  • After measuring the beat, we swapped out the SHP–150 on the RF input of the mixer for an SHP–100. The beat is currently at 120 MHz, and with the SHP–150 we were throwing away something like 60% of our power. Attenuation from the SHP–100 appears negligible when viewed on a scope.

The beat spectrum is attached, along with the expected coating Brownian noise estimate. I will post the estimates of the PDH and RIN contributions later.

Attachment 1: beat_2013-11-04.pdf
beat_2013-11-04.pdf
  1381   Tue Nov 5 01:11:15 2013 taraDailyProgressElectronics EquipmentTTFSS

Quote:

 

 Do you guys have a plot that shows the required loop gain and the achieveable loop gain with this TTFSS on the same plot?

 Not yet, we will add this later. but we measured the noise at error point before and it is well below the estimated coating noise.

 

Plan for this week

Mon: (See Evan's entry for more detail)

  • minimize RFAM,
  • bring back the beat signal, optimized gain setup. A lot of improvement around 10 - 50 Hz.
  • measured error noise from each PDH loop, with slope of the error signal to calibrate it to frequency noise
  • measure RIN

Tue:

  • measure photothermal coupling from RCAV (may remeasure ACAV)
  • measure cavity pole, to check the cavities' finesses
  • check calibration on PLL

Wed

  • replace the broken air spring with a new one.
  • fix harmonic lines in the beat
  • update the calculation for the noise budget ( electronic noise, etc)
  • Turn on ISS

Thur

  • beat measurement
  1380   Sun Nov 3 19:37:32 2013 ranaDailyProgressElectronics EquipmentTTFSS

 

 Do you guys have a plot that shows the required loop gain and the achieveable loop gain with this TTFSS on the same plot?

  1379   Fri Nov 1 00:22:40 2013 taraDailyProgressopticmore optimization

I'm putting EOAM back on ACAV path. The setup is roughly optimized.

(14.75 MHz) EOM , EOAM, quarter waveplate and PBS in ACAV path are put back together. I used a half waveplate in front of the EOM to adjust the beam to S- polarization. Right now all the polarizations optimization (to all EOMs, both ACAV/RCAV path) are adjusted to S-polarization with respect to the table. We may have to fine tune it later to match the E field in the EOMs.  The EOAM setup is optimized. With +/-4 V, the output power can be adjusted to 1mW +/- 0.09 mW (+/- 9%). The performance is comparable to RCAV EOAM. (10%) . I have not add another half waveplate before the EOAM yet. We can add it back later if we need to adjust the input polariztion to the EOAM.

I checked scattered light in the area between PMC and ACAV.  There is a reflection from EOAM back to EOM, but I cannot really block it with an iris. It probably bounces of the case of the EOM or going back to the crystal. Anyway I'll block the beam around this path later.

I have not aligned the beam to the cavity yet, since the temperature was changing because I removed the insulation  caps to patch them with black out material.

 I put black out material (R @1064 ~0.4-0.6%)on the vac tank insulation caps to minimize any possible scattered light source inside the tank that might leak out.  It also keep the surface cleaner from all the foam dust.

foam1.jpg

foam2.jpg

  1378   Thu Oct 31 18:10:07 2013 EvanDailyProgressElectronics EquipmentTTFSS

Tara and I spent some time looking at the TTFSS boards. The offset issues appear to be caused by bad choice of offset knob on the TTFSS interface boards. Previously, Tara and I had used the offset knob to null RFAM-induced offsets in the north PDH signal. The current thinking is that when Tara re-optimized the electro-optic elements on the north path, the RFAM-induced offset changed and was therefore no longer nulled.

We have now returned the offset knobs to their optimal values as follows: we set each TTFSS to use the TEST SMA input rather than the LO/PD + mixer input, we applied a 50 Ω terminator to this input, and we then watched TP4 (after the common VGA) on a scope while adjusting the offset knobs. The optimal knob positions are 526 for the south TTFSS and 506 for the north TTFSS. Varying the common gain causes the DC offset on TP4 to change only slightly (it stays within ±5 mV for both north and south). Varying the fast gain causes the DC offset on TP17 (after the fast VGA) to vary as well; on south, this also appears to stay within ±5 mV of zero, but on north it is as high as 20 mV when the common and fast gains are turned all the way up. However, since these VGAs are each +30 dB at maximum gain, this means that the offset referred to TTFSS OUT1 is more like 20 µV, which is negligible compared to an error signal that is something like 1 Vpp.

  1377   Thu Oct 31 00:02:17 2013 taraDailyProgressElectronics EquipmentTTFSS

Evan found that when common gain is changed, DC offset also changes as well. I'm still looking into the problem.

 

D040105C.png

a part of schematic, the driving signal was sent in through test port (the switch was flipped from off to test), so the signal came through PD line in this page.

 We still cannot lock RCAV with TTFSS, so I'm checking the box 2009007 (#7).

  • The modifications I did two days ago were 1) adding a push switch for gain reduction and 2) replacing one resistor. The rests were changed before we got the TTFSS.
  • I checked the TF between TP1 and TP5, it works as it should be (20log( 390/100) ~ 12 dB). So this stage does not have any problem. I checked both TF and time domain signals. So the modifications are ok.
  •  Note, when I measured the TF of TP1/TP4 or TP5/TP4, the signals oscillated and became very noisy. I don't understand why, but this problem disappeared when I used TP4 and out2. Both boxes (#5,#7) have this same problem.

Common Gain - DC offset problem

  • When Common gain is increased (CG signal to U2A chip), there is an offset observed in TP4 which is after the variable gain stage. Both boxes behave similarly. <- This surprises me, as we haven't seen this (or haven't noticed this) before.
  • I checked if the offset varied with the input drive or not. I changed the input from 20mV to 40mV, with constant gain = 1000(25dB). The behavior is nonlinear (see the plot below). I checked this only on box#5.

offset.png

DC offset vs input drive. DC offset is calculated from (Vmax + Vmin) /2  from a sinusoidal signal input. The signal was taken from TP4. The behavior is very non linear and it is impossible to make a table for an appropriate offset level vs common gain setting.

 What to do next?

  • This seems to be an important clue about why the loop behaves badly when common gain is increased. From today test, both boxes behave the same, so I think it might be the chips' problem.
  • Fortunately, we can use the offset adjustment(OS)
  • to cancel the offset introduced by the common gain, but we might need to add a port some where (we might be able use fast mon channels during laser scan).  So that when we increase the gain, we can adjust the offset accordingly.
  • Box#7 that I modified should be working. I don't know why I could not lock the laser before, more checking has to be done.

 

  1376   Wed Oct 30 01:56:38 2013 taraDailyProgressoptictable work

I'm optimizing the setup, and clearing the table a little bit.

  • Self homodyne setup in ACAV path is removed. This is from Erica's setup and it is not used. The input part is left, since I might use it for fiber distribution system
  • optics on RCAV path, all polarization are optimized. This includes, the input and output polarization for EOAM, and quarter wave plate before the periscope. The input polarization for sideband EOM is left intact after the last adjustment, and it should be good. With+/- 4V input, I can change the power by +/-10%, (1.0 +-0.1 mW is the current setup). For Evan: Do not touch anything before discussing with me!!!
  • I replaced a new PBS for PDH locking in RCAV path. The old one is bad. The surface between the prisms is milky, see the pictures below for comparison. There is also beams from multiple reflection within the cube. The new one is much better. There is no ghost beam anymore.
  • I blocked all the scattered light I could find in RCAV path with Irises and beam dumps. For ACAV, I just blocked the scattered lights from the laser to the PMC. I will finish the whole setup later.
  • I rechecked the height of the beam through EOMs/EOAMs. Since it is a little tricky to center the beam through the openings. The EOMs in RCAV path are all checked. For ACAV, only those between the laser and the PMC are checked(BB for phase locking and 21.5 for PMC sideband). The 14.75Mhz sideband and EOAM will be done later. The EOAM and wave plates are removed temporarily.
  • I modified the TTFSS for RCAV to have a gain reduction switch to help locking the laser. I tried to lock RCAV, but I cannot turn up the gain. I'm not sure what I did wrong but this has to be investigated.

To do lists

  • put optics back in ACAV path and optimize them (alignment + polarization).
  • fix RCAV TTFSS . Check by measuring the TF of the modified stage/ scanning laser + checking error signal

oldPBS.jpg

above: old PBS, bad inter surface can be seen.

newPBS.jpg

above: new PBS: all surfaces are clear

  1375   Mon Oct 28 22:55:11 2013 EvanDailyProgressBEATBeat measurement with ISS

I've taken Tara's farsi.m and changed the values of finesse F and absorption α in order to fit the magnitude of the TF measurement in PSL:1368. I've chosen 7500 for the finesse and 5 ppm for the absorption, although for this calculation they are degenerate (entering into the TF as F/α).

Using this, I've taken the RIN measurements from Friday and used them to estimate the induced frequency fluctuation in the beat readout, assuming a transmitted power of 1 mW from each cavity.

In the case when the ISS is off, the estimated effect of RIN on the current beat is significant only below 10 Hz. When the ISS is on, the RIN is insignificant over the entire measurement range. This perhaps explains the observed reduction in the beat PSD below 10 Hz when the ISS is on.

Attachment 1: rin_to_beat.pdf
rin_to_beat.pdf
Attachment 2: expected_rin-to-beat.pdf
expected_rin-to-beat.pdf
Attachment 3: expected_rin-to-beat_iss.pdf
expected_rin-to-beat_iss.pdf
  1374   Sun Oct 27 20:12:25 2013 taraNotesoptic photothermal noise in AlGaAs

I revised the calculation for photo-thermal noise in AlGaAs coatings, the photo thermal noise should not be a limiting source.

==review==

photothermal noise arises from the fluctuation in the absorbed laser power (RIN + shot noise, mostly from RIN) on the mirror. The absorbed power heats up the coatings and the mirror. The expansion coefficient and refractive coefficients  convert thermal change into phase change in the reflected beam which is the same effect as the change of the position of the mirror surface.

Farsi etal 2012, calculate the displacement noise from the effect. The methods are

  • Solving heat equation to get temperature profile in the mirror.
  • Use elastic equation to calculate the displacement noise due to the temperature change (thermoelastic)
  • For TR, the effect is estimated from effective beta (from QWL stack) and the temperature at the surface ,as most of the TR effect comes from only the first few layers

When they solve the heat equation, the assume that all the heat is absorbed on the surface of the mirror. This assumption is ok for their case ( SiO2/Ta2O5) with Ta2O5 at the top surface, all QWL, as 74% of the power is absorbed in the first four layers (with the assumption that the absorbed power is proportional to the intensity of the beam, and all absorption in both materials are similar).

However, for AlGaAs coatings with (nH/nL) = (3.48/2.977) The E field goes in the coatings more that it does in SiO2/Ta2O5, see the previous entry. So we might want to look deeper in the calculation and make sure that photo thermal noise will not be a dominating noise source.

==calculation and a hand waving argument==

 The plot below shows the intensity of the beam in AlGaAs Coatings, opt4, and the estimated intensity that decreases with exponential square A exp(-z^2/z0^2). X axis is plotted in nm (distance from surface into coatings). The thickness of opt4 is about 4500 nm. To simplify the problem, I use the exponential decay function as the heat source in the diff equation. But I have not been able to solve this differential equation yet. Finding particular solution is impossible.  So I tried to solve it numerically with newton's method, see PSL:284. But the solution does not converge. I'm trying green function approach, but i'm still in the middle of it.

Int_cotings.png

 

However, the coatings optimized for TO noise should still be working. Evans etal 2008 discuss about how the cancellation works because the thermal length is longer than the coating thickness. The calculation (TE and TR)  treat that the temperature is coherent in all the coatings ( they also do the thick coatings correction where the heat is not all coherent, and the cancellation starts to fail at several kHz). So the clue here is that the cancellation works if the heat (temperature) in the coatings change coherently.

For photothermal calculation, if we follow the assumption that all heat is absorbed at the surface (as in Farsi etal), we get the result as shown in psl:1298, where most of the effect comes from substrate TE . In reality, where heat is absorbed inside the coatings as shown in the above plot, heat distribution in the coatings will be even more coherent, and the effect from TE and TR should be able to cancel each other better. Plus, higher thermal conductivity of AlGaAs will help distribute the heat through the coatings better.

This means that  the whole coatings should see the temperature change more coherently, thus allowing the TO cancellation in the coatings to work. The assumption that heat is absorbed on the surface should put us on an upper limit of the photothermal noise.

This means that photothermal noise in the optimized coatings should be small and will not be a dominating source for the measurement.

 

Attachment 2: Int_cotings.fig
  1373   Sat Oct 26 15:35:42 2013 EvanDailyProgressBEATBeat measurement with ISS

Summary: No good so far. Engaging the ISS seems to have basically zero effect on the beat. The beat overall looks worse than it did a month ago, and the shape seems to mimic the shape of the north cavity RIN. More optimization of the north EOAM is necessary.

Details: Having set up the north EOAM on Thursday (PSL:1372), I spent most of yesterday trying to get a RIN-suppressed beat measurement.

The continual drift of the laser frequency control signals was irritating, so I spent some time getting the slow digital PID controls for the lasers back up and running. At first only KP seemed to have no effect on the laser control signals; it turns out this is because the PID Perl scripts that run on the Sun machine rescale the KI and KD coefficients by a timestep variable, which had been set to zero. I've set it to 1. I've chosen KP = KD = 0 and KI = 0.0002 (with appropriate choice of sign for the two loops). The system is probably overdamped, but it manages to integrate the control signals down to zero in a resonable amount of time (<30 s) and I don't think it's a high priority to optimize it right now.

The south PDH error signal has noticeable 250 kHz oscillations which get worse as the common TTFSS gain is increased. The north PDH error signal is much quieter. Are we perhaps hitting a mechanical resonance of the EOM crystal? Or (dare I say it) do we have the wrong sign for the common path of the PDH loop?

I took out the hand-soldered integrating board that I built for the ISS loops; it was railing too often. The ISS setup for each path is now as follows: each ISS PD goes into the A input of an SR560, and a programmable voltage reference (Calibrators Inc. DVC–350A) goes into the B input. The voltage is chosen to match the dc voltage from the ISS PD. The SR560 is dc coupled and set to take the difference A − B. The gain is set to 5×103 V/V, with a single-pole low-pass at 1 kHz. The output from the SR560 is fed into the EOAM.

The suppressed and unsuppressed RIN measurements are given in the first two plots. Evidently, these simple ISS loops are able to suppress the RIN by a factor of 50 or so. Also, the north RIN is much worse than the south RIN, and the hump from 100 Hz to 10 kHz is reminiscent of a poorly aligned EOAM (as seen in PSL:1311, for example). So I'd like to spend some more time fiddling with the north EOAM to see if I can improve the RIN suppression. Alternatively, perhaps we are suffering because the north path has no PMC to stabilize the pointing into the EOM, EOAM, etc.

Anyway, I pressed ahead and looked at the beat. To convince myself of the repeatability of the setup, I took a measurement with the ISS loops on, then a measurement with the ISS loops off, and then a measurement with the ISS loops on again. The result is given in the third plot. Below a few hertz, the ISS may have a positive effect. Above this, there is either no effect or a small worsening effect.

Note that the shape of the beat follows the shape of the north cavity RIN. I think we should spend a little time noise hunting and optimizing on the north path to see if we can make this go away. Note also that the beat is worse than it was back in September (PSL:1321). Two immediate culprits that I can think of are (a) the installation of the EOAM or (b) the fact that the vacuum can is no longer floated. But it could just as well be that there's something else (e.g., PDH offsets) that I neglected to optimize.

Attachment 1: north_iss.pdf
north_iss.pdf
Attachment 2: south_iss.pdf
south_iss.pdf
Attachment 3: beat.pdf
beat.pdf
  1372   Thu Oct 24 16:44:43 2013 EvanDailyProgressISSNorth EOAM: working

Summary: The EOAM is back from Newport and it looks like it's working. With 3 mW incident on the EOAM, we get 1 mW after the PBS. I applied a range of DC voltages between −4 V and +4 V and measured the output power. The effect is linear, with a slope whose magnitude is 17 μW/V. From the manual, we expect a slope of (π / 2) (3 mW / 300 V) = 16 μW/V, so we're pretty much spot on (assuming a half-wave voltage of 300 V).

Details: From upstream to downstream, the EOAM setup consists of a HWP, the 4104 amplitude modulator, a QWP, and finally a PBS.

  • With the EOAM and QWP removed, I adjusted the HWP in order to maximize transmission through the PBS. The original angle was 98.5°, and I turned it to 94.0°. The power onto the PBS was 1.19 mW, the PBS transmitted power was 1.11 mW, and the PBS reflected power was 47 µW.
  • I added the QWP back in and rotated it to maximize transmission through the PBS. (This was my interpretation of Newport's instruction to include "a quarter waveplate oriented vertical to the modulator housing".) The power onto the PBS was 1.19 mW, the PBS transmitted power was 1.12 mW, and the PBS reflected power was 35 µW. The QWP angle was 100°.
  • I added the EOAM and drove it with a 10 Hz, 6 Vpp sine wave. At this point I found that my choice of QWP angle (100°) was incorrect, so I tried two other angles: 145°, which should put the QWP axes at ±45° relative to vertical, and 132°, which I found is close to equalizing the transmitted and reflected powers from the PBS (this should in principle coincide the correct operating point for the EOAM). For each of these angles, I measured the reflected and transmitted PBS powers, and I watched the DC output of the north refl PD in order to get the depth of the intensity modulation. These numbers are given in the table below.
QWP angle (deg.) 100 145 132
Refl PD DC mean (mV) 372±4 139±2 197±3
Refl PD DC peak-to-peak (mV) 4.0±0.5 13.6±0.4 12.8±0.4
Modulation depth (%) 0.54±0.07 4.89±0.16 3.25±0.11
PBS transmission (µW) 32±2 409±2 570±2
PBS reflection (µW) 1060±10 677±2 532±2

In the end, I decided to use 145° as the QWP angle since it provided the greatest modulation depth of the three angles that I tried. Additionally, it jives with my understanding of the EOAM setup; namely, that the 4104 by itself acts as a voltage-controlled waveplate that (a) has its axes located at ±45° and (b) has zero retardance in the absence of applied voltage. Therefore, to bias it to have λ/4 retardance, one should add a QWP with its axes at ±45°.

After I took these measurements, I then rotated the HWP after the north EOM from 192° to 193° in order to get 3.0 mW incident on the EOAM setup. I then took the calibration data plotted below using the handheld voltage reference and the ThorLabs power meter.

Attachment 1: north_eoam.pdf
north_eoam.pdf
  1371   Wed Oct 23 15:16:25 2013 ChloeDailyProgressECDLFree Running Noise Experiment

 

Note: last week I picked up the modified diffraction grating mount. I forgot to bring it in today but I'll put it back in the ATF lab on Thursday. 

I've spent the last week reading a few papers Tara sent me about mode matching/old elog entries by various people. I couldn't find Tara around the lab today, so I'll try and talk to him this week to figure out exactly what I'm doing. I'm still a bit confused about how to do the setup, although I've been starting to sketch what I'm planning on doing. I'm also messing around with a few mode matching programs to help me plan my setup. 

I sent up the power supply for the PD and confirmed it works. I'm going to try to talk to Tara tomorrow or Friday so I know what I need to do in the next week. My midterms are starting so I may have a hard time being around the lab much until afterwards. 

  1370   Tue Oct 22 04:34:12 2013 taraDailyProgressSeismicnew table legs installed

After installing the table legs, I have been trying to measure the beat. However, there is an unknown scattered light noise up to 400 Hz. I'm still trying to fix that.

  Here are some bullets about what happened, I'll add the details later.

  • Extra noise that looks like scattered light goes up to 400Hz ( was around 100 Hz before, not from floating the table)
  • One of the air spring supporting the vac tank has a leak. But it is unlikely to be the source of the extra noise mentioned above.
  • The finesse of the cavities may be less than the designed value (10 000) because of the not so clean isoprop I used on the mirrors.

Note: check if the beams in the tank is blocked by wires or not.

  1369   Mon Oct 21 14:04:04 2013 ranaDailyProgressBEATNew intensity-to-frequency TF

 Even though it shouldn't be there, there is probably a coupling directly from RIN to the reflected PDH signal on each cavity. So you should measure all 3 terms (RIN 2 PDF for each cavtiy, and RIN 2 Beat).

  1368   Thu Oct 17 21:50:28 2013 EvanDailyProgressBEATNew intensity-to-frequency TF

[Tara, Evan]

We took another RIN-to-beat transfer function, as in PSL:1316. This time we've directly measured the conversion factor between power transmitted through the south cavity and voltage put out by the transmission PD. To do this, Tara purposefully misaligned the alignment into the cavity in order to get several different transmitted powers. For each misalignment, we measured the power immediately after the vacuum can using the power meter as well as the voltage out of the south PDA10CS with its gain at 30 dB. The result is (1.78 ± 0.05) V/W, and the fit is shown in the first attachment.

The second attachment shows the transfer function. I've applied the PD conversion factor mentioned above, as well as the Marconi actuation conversion factor (710 Hz/V for the 1 kHz FM setting). The red and green traces were taken with the table not floated, and the blue traces were taken with the table floated (we originally took all the traces with the table not floated, but the SR785 decided to write an empty data file and we didn't realize it until after we floated the table).

Also, I think I must have applied the wrong calibration (7.1 kHz/V) in PSL:1316; at low frequencies, the TF there is almost exactly a factor of 10 higher than the TF here.

Attachment 1: trans_pd_cal.pdf
trans_pd_cal.pdf
Attachment 2: rin_to_beat.pdf
rin_to_beat.pdf
Attachment 3: ctn_rin_beat_2013-10-17.zip
  1367   Mon Oct 14 21:02:00 2013 taraNotesopticcoating optimization for AlGaAs:variation in x

I checked the dependent of coatings properties with the uncertainty in x (amount of Al in Al_x Ga_(1-x) As). The effect is already within the uncertainties in materials parameters we did before and will not be a problem.

G. Cole told us about the variations in Al contents in the coatings. Right now the values are 92% +/- 0.6%. 

(92.10, 91.43, 91.34, 91.57, 92.73, 92.67).  Although the deviation is small, the Al content does not always hit 92%, but 92+/- sigma%. So I decided to check the effect of x on the optimization.

The materials properties that change with x are heat capacity, alpha, beta, heat conductivity and n. The values of those as functions of x can be found on ioffee  except n. So I looked through a couple of sources ( rpi, sadao)  to get n as a function of x, (Note: E0 and D0 are in eV, they have to be converted to Joules when you calculate chi and chi_so).  GaAs (nH) has a well defined value ~ 3.48+-0.001, nL has a bit more uncertainty, but it is within the approximated standard deviation of 0.03 . The table below has numbers from the sources. For RPI, I use linear approximation to get nL for x = 0.92 @ 1064nm.

source nL(x=0.92) nH
G.Cole 2.977 3.48
RPI 3.00 3.48
Sadao 2.989 3.49
     

The dependent of n on x is about -0.578 *dx. The numbers from RPI and Sadao are about the same. This means that for the error of 0.6% in Al. nL can change by 0.578*0.006 = 0.0035. The number is almost a factor of ten smaller than the standard deviation of nL and nH I used in previous calculation (0.03). For examples,

  • x = 0.914, nL = 2.993,
  • x=0.92,     nL = 2.989
  • x=0.926    nL = 2.986  (From Sadao's fit)

This means that the uncertainty in nL/nH (+/- 0.03) we used are much larger than the effect coming from uncertainty in x. This is true for other parameters as well.

  1366   Mon Oct 14 17:38:10 2013 ChloeDailyProgressECDLFree Running Noise Experiment

 

I got the modified grating design into the Caltech machine shop; this part should be done by tomorrow. We decided to use 2 vertically placed 1/4-80 holes which will have adjustment screws. This will allow for tilt adjustment. 

I found a mixer and a splitter in the CTN lab plus the appropriate adapters to use. I'm still working out how the cable length difference will affect the sensitivity of our measurement. 

I have the PD removed from the Gyro table along with the lenses that were used to focus the beam to go into the PD. This was more difficult than I expected to remove these pieces since I'm short and didn't want to disturb the other setup. We are still need several things:

  • Tara helped me find a power supply I can use for the PD
  • stable laser light from fiber optic from CTN lab: Tara said the laser in the CTN lab is working but I need to work out the mode matching. I will read up on this/Erica's old elog posts. 

I'll go pick up the modified grating mount from the Caltech machine shop tomorrow so that I can wash it tomorrow afternoon (I don't have much time tomorrow) and do more on Wednesday. 

  1365   Fri Oct 11 15:23:54 2013 taraNotesopticcoating optimization for AlGaAs:electric field in coating layer

Electric field in coating layer is calculated. This will be used in loss calculation in AlGaAs coatings.

 

  • In each coating layer, there are two E waves, transmitted and reflected  waves. The two interfere and become an effective field.
  • The averaged electric field will depend only on the transmitted beam inside each layer, see the calculation.
  • The effective transmissivity can be calculated, for coatings with N layers between air and substrate, there will be an N+1 vector representing the effective transmission, called tbar in the code. This tbar(n) is the transmissivity in the nth layer, similar to rbar in Evans etal calculation.
  • The ratio of E field/ E input in nth layer will be tbar(1)*tbar(2)*...tbar(n)
  •  |E field/ E input |^2 of the final transmitted beam is the transmission of the coatings.  The numbers from this calculation agrees to the calculation from before.

==supplementary information==

1) average E field in layer is the transmitted E field in the layer.

avgE.jpg

 I attached a short matlab file for a simulation of the combined field. Ein in each layer will be the transmitted beam through the layers. For a value of r close to 1, we get a standing wave. Try changing the value of r in test_refl.m to see the effect

 

2) Calculation for the transmitted field in each layer

transE.jpg

I borrow the notation from Evns etal paper (rbar), the calculation code multidiel_rt.m is attached below. Note: the final transmission calculated in the code is the transmission from the coating to the substrate. To calculate the transmission to the air, multiply the last transmission by 2*n_sub/(n_sub + n_air) which is the transmission from sub to air. Since the thickness of the substrate is not known with the exact number, it will not be exact to the transmision calculated in GWINC or Matt A's code (which do not take the sub-air surface into account), but they will be close, because the reflected beam in the last interface will be small compare to those in the coatings.

 

==result==

Efield.png

The penetration of E field for QWL and different optimized coatings are shown here. The transmissions in the legend are calculated from MattA./GWINC and the values in the parenthesis are calculated from multidiel_rt.m which include the effect from the substrate-air surface. This makes the values in the parenthesis smaller (as more is reflected back and less is transmitted).

Attachment 3: test_refl.m.zip
Attachment 4: multidiel_rt.m.zip
Attachment 6: Efield.fig
  1364   Thu Oct 10 16:59:59 2013 ChloeDailyProgressECDLFree Running Noise Experiment

 Today, I met with Tara and discussed the delay line, which will be used to tune the wavelength of the bare laser diode and measure the free running noise of the ECDL. I will write up notes of what we talked about and how the delay line will work and post these soon, along with a list of items I will need. 

Collecting materials for delay line: I will be using an RF photodiode and the series of lenses from the Gyro setup, which Evan is not using right now. I'm in the process of disassembling and reassembling this setup on the table that I'm using. Evan said the mode matching was already messed up, so I will be working on focusing the beam. I will be using NPRO light from the CTN lab via a fiber cable. 

Tara helped me modify the piece I found last time, which goes under the collimating lens mount to fix the optical height. I learned how to tap a hole in the metal. I moved the ECDL setup and got the current driver back up and running, and was able to focus the beam using the collimating lens we purchased. The setup so far is attached. 

Photo_on_10-10-13_at_4.49_PM.jpg

Tara and I decided on a modification to make to the grating mount which will allow for us to make vertical tilt adjustments (we will have 2 holes with adjustment screws, not one). I am going to draw this in Solidworks so that I can get it machined tomorrow at the Caltech machine shop. 

  1363   Thu Oct 10 01:59:24 2013 taraNotesopticcoating optimization for AlGaAs:error analysis

I recalculated the coatings properties, with the values of nH and nL to be 3.48 and 2.977. Note about each optimization is included here. Transmission plots are added in google spread sheet. I'll finish the calculation for E field in each layer soon.

Note about each optimized coating version: different versions were obtained from different cost functions, and different number of layers.

opt1

  • 55 Layers
  • T = 210 ppm
  • TO noise and transmission is too sensitive to the change in nH and nL
  • 1/4 cap of nH. I did not fix the cap thickness for other coatings. Since there is no reason to keep the thickness of the cap constant.
  • TO noise and transmission of this one changes a lot with uncertainty in nH/nL

 

opt3

  • 57 Laayers
  • T = 150 ppm
  • Transmission is still too sensitive to the change in nH and nL
  • TO noise/ transmission is less susceptible to change in nH/nL.
  • First layer is 0.1 lambda thick (~285 nm) I'm not sure if this will be a problem for a cap or not.

 

opt4

  • 57 Layers
  • T = 150 ppm
  • TO noise and Transmission are less sensitive to nH and nL
  • less amount of nL material, should be less sensitive to error in thickness control

 

opt5

  • 59Layers
  • T= 144 ppm
  • TO noise and Transmission are less sensitive to nH and nL
  • reflected phase is more sensitive compared to opt4
  • use less nL material
  • 0.1 lambda thick

Judging from TO noise level, Transmission and reflected phase, I think opt4 is the best choice for us. The structure consist of thick nH layers and thin nL layers. This is good for us in terms of thickness control.

 

  1362   Wed Oct 9 16:12:39 2013 ChloeDailyProgressECDLTEC Construction

Tara mentioned that the TEC may not work as well without some of the silicone thermal paste so I added some and returned the PID gains. Sure enough, this helped the temperature stabilize (and now I know if it stops working to clean out the Peltier element with isopropanol). 

I emailed DMass about finding a PCB for building the low noise current controller. (I was supposed to do this last week but it slipped my mind)

I moved the laser diode and socket to the actual laser diode mount (from the Michelson setup used in August). Since the laser diode extrudes, we do not have the problem I mentioned last time with the base plate since the hole for mounting the collimating lens is now close enough that we should have enough adjustability to focus the beam. I searched the ATF lab and found a piece of metal about 3mm high, which will fix the difference in height of the diode mount and the collimating lens mount. However, this piece needs to be trimmed down, which I will try to discuss with Tara. Not sure if we have those capabilities here or if I need to take it to the machine shop?

I wasn't able to find Tara today but I need to talk to him about:

  • grating mount modifications to allow for fine adjustment of tilt when gluing the grating... I have no idea what to design for this since every idea seems very complicated. Maybe we can try it and see how much this is necessary before proceeding? Nothing in the literature has suggestions about how to mount this; most people seem to have just glued it on. I was thinking we add a second adjustment screw to the grating mount, but I'd like to discuss with Tara about this
  • finding materials to test wavelength... I searched for a RF photodiode but was unable to find anything with a bandwidth larger than 50 MHz in the ATF lab. I will try to look again for this. I also need to talk to Tara about how we plan to get stable laser light to interfere with the ECDL beam.
  • for a few weeks down the line...how do I design a delay line? How does it work? I will try to read up on this on my own. 

Finally, I'm still working on editing my SURF paper. I'm new to LaTex so it's taking me awhile to perform the edits Rana suggested. 

  1361   Tue Oct 8 23:01:56 2013 EvanDailyProgressBEATAttempts at new beat measurement

[Tara, Evan]

Having successfully floated the table yesterday, we attempted a new beat measurement in the hopes that the large shelf below 100 Hz had disappeared. Unfortunately, this appears to not be the case. Additionally, many of our signals are plagued by unusually large, slow drifts. We're hoping that they're just thermal transients caused by all the work on the table over the past 12 hours, and that by tomorrow things will have settled down. We'll see if that's the case.

Anyway, we did the following things today:

  • We reconnected cables that come in from off the table and go onto cameras, PDs, etc., paying special attention to strain relief and vibration isolation since the table now floats.
  • We redid the alignment to recover ~90% visibility; this required only touching the periscope mirrors (somewhat surprising considering what we subjected the table to in order to switch out the legs).
  • We got the cavity PDH loops up and running again. The control signals show unusually large drifts. We also noticed this while sweeping the lasers to align the cavities; the resonance for north in particular would wander out of the sweep range every 30 seconds even though the laser was being driven at 10 Vpp from an SRS function generator.
  • We spent some time trying to null (what we assume is) RFAM-induced offset in the PDH error signals. We did this by adjusting the HWPs before each cavity EOM and nulling the offset on TTFSS common OUT1. The south cavity already had a small offset, so no adjustment was required. On the north cavity, there was a noticeable offset (~20 mV, compared to an error signal pk-pk of 220 mV), so Tara nulled it. We then found that we could get a stable lock with the laser PZT actuator alone, and that adding the EOM actuator caused the loop to oscillate (almost as if the EOM actuator was driven with the wrong sign). So we looked at the error signal again, and unexpectedly found another ~20 mV offset. Tara nulled it again and this time the lock was stable; in fact, we were able to get the common and fast gain knobs up to 1000 and 1000 (compared to 800 and 800 earlier in the day). No idea what the problem is here; possibly it drifted between the successive adjustments.
  • We looked at the beat. It appears to be not much better than with the table unfloated.
  • We took a measurement of RIN-induced disturbance in the beat by driving the south EOAM with a sine from the SR785 and taking the TF that takes transmission PD intensity to beat fluctuation. Unfortunately, this measurement is not consistent (in magnitude or phase) between successive sweeps. It seems to be due (at least in part) to DC drift in the beat.
  • We tried turning on the crude south ISS, but it made the beat more noisy.

 

ELOG V3.1.3-